# Multi-Agent Diagnostics for Robustness via Illuminated Diversity

### v1

@@ 5. Results and Discussion
In our analysis of targeting TiZero on GRF, we closely examine the performance of MADRID and baselines. Figure 3(a) displays the average estimated regret values for all $160$ cells within the MAP-Elites grid across the entire collection of reference policies. Here, MADRID outperforms both baseline methods. The random baseline exhibits a negative value close to 00, as TiZero proves to be a stronger policy than all the reference policies on entirely random game levels. On the other hand, the targeted baseline performs well, closely resembling MADRID’s performance at the early stages of iterations. However, as the iterations continue, it lags behind due to its failure to capitalise on previously identified high-regret levels that serve as stepping stones for next iterations.
In Figure 3(b), the variation in estimated final regret scores across the reference policies is illustrated. Here, the regret increases as we move to higher-ranked agents. The heuristic bots display regret levels that are on par with the intermediate checkpoints of TiZero.
As we approximate the regret using the difference between cross-play (XP) and self-play (SP) between reference and TiZero policies (see Equations 1 and 2), a regret estimate of $1$ for an adversarial level $\theta$ can be achieved in two situations. First, the reference policy scores against TiZero in XP, while TiZero cannot score in its SP. Second, TiZero concedes a goal in SP in $\theta$. Intriguingly, our findings reveal that for around 90% of the adversarial levels generated by MADRID, a nominally weaker reference policy outperforms TiZero. This emphasises MADRID’s capability in exposing adversarial levels where even state-of-the-art policies be prone to missteps.
@@
Figure 3(c) and Figure 3(d) illustrate the estimated rate of goals scored against TiZero by the reference policies on adversarial levels produced by MADRID and baselines. We can see that in approximately 70% of the time across all reference policies, the reference policy scored a goal against TiZero in a short period of time. [^6] The levels last only $128$ environment steps, which is a short episode compared to the $3000$ steps for the full game. It should be noted that within the remaining 30%, the majority of instances resulted in no goals due to the nondeterministic dynamics of the environment.
Figure 5 highlights the difference in performance for selected reference policies. Notably, the higher-rank checkpoints of TiZero, saved at the later stages of its training, can be used to identify more severe vulnerabilities, as measured using the regret estimate.
Figure 6 shows the evolution of MADRID’s archive for a specific reference policy, illustrating its search process over time. Initially, the grid is sparsely filled with low-regret levels. However, as iterations progress, MADRID generates high-regret levels that progressively populate the entire grid. This shows that MADRID can discover high-regret levels anywhere on the football field. On average, we notice that higher-level scenarios tend to be located towards the positive $x$ coordinates. These correspond to situations where the ball is close to the opponent’s goal from the perspective reference policy. While most regret scores tend to have uniform values around in similar positions on the field, in Figure 5(d) the grid also includes an adversarial level with estimated regret of $1.75$. This indicates that MADRID found a level where the reference policy scores against TiZero in XP, while TiZero concedes a goal in SP.
@@ 5.1. Qualitative Analysis
Next we conduct a qualitative analysis of the adversarial levels identified by MADRID on GRF by visualising the highest ranking levels in the archive across all reference policies. We provide a selection of these examples below, with a comprehensive list available in Appendix A. Full videos of all identified vulnerabilities can be found at https://sites.google.com/view/madrid-marl.
@@
(b) The receiving player is clearly in offside, thus a freekick is awarded to the opponents team.
(c) Reference policy does not pass to offside player and directly runs towards the goal to score.
@@
Despite its strong performance under standard evaluations, TiZero frequently falls victim to erroneously passing the ball to players unmistakably in offside positions, as shown in Figure 7 This observations highlights TiZero’s lack of a deep understanding of the rules of the game. In contrast, the reference policies abstain from passing the ball to offside players, resulting in successful scoring outcomes. [^7] A player is offside when it is in the opponents’ half and any part of their body is closer to the opponents’ goal line than both the ball and the second-last opponent. Usually one of the two opponents is the goalkeeper. When this happens a free kick is awarded to the opponent’s team.
@@
Figure 9. Adversarial example of a slow running opponent. Three TiZero-controlled defenders are not able to stop a simple slow running opponent, who walks past them and scores.
@@
(c) Reference policy passes the ball to a better positioned player who scores.
@@
Perhaps the most glaring adversarial behaviour discovered are instances where TiZero agents inexplicably shoot towards their own goal, resulting in unforced own goals (See Figure 8). In contrast, when starting from identical in-game positions, the reference policies manage to counterattack effectively, often resulting in successful scoring endeavors.
Slow running opponents
The TiZero agents always choose to sprint throughout the episode. However, this makes them weak on defense against opponents who move slower with the ball. Instead of trying to tackle and take the ball, TiZero’s main defensive strategy is to try and block opponents. Opponents can take advantage of this by using deceptive moves, especially when moving slowly, making it hard for TiZero’s defenders to stop them. This is illustrated in Figure 9.
@@
A notable shortcoming in TiZero’s policy, when compared to the built-in heuristic, is its reluctance to pass the ball to teammates who are in more favorable positions and have a higher likelihood of scoring, as illustrated in Figure 11. In contrast, heuristic bots—whether easy, medium, or hard—demonstrate a consistent pattern of passing to optimally positioned players, enhancing their goal-scoring opportunities. This effective passing strategy seems unfamiliar to TiZero, causing it difficulty in overcoming a successful defense.
@@ 6. Related Work
Quality Diversity (QD) is a category of open-ended learning methods aimed at discovering a collection of solutions that are both highly diverse and performant(Lehman and Stanley, 2011; Cully and Demiris, 2018). Two commonly used QD algorithms are Novelty Search with Local Competition (NSLC, Lehman and Stanley, 2011) and MAP-Elites (Mouret and Clune, 2015; Cully et al., 2015). These two approaches differ in the way they structure the archive; novelty search completely forgoes a grid and opts instead for growing an unstructured archive that dynamically expands, while MAP-Elites adopts a static mapping approach. Although MADRID leverages MAP-Elites as its diversity mechanism, it can be adapted to use NSLC. One of the most effective versions of MAP-Elites is CMA-ME (Fontaine et al., 2020). CMA-ME combines MAP-Elites with the evolutionary optimization algorithm Covariance Matrix Adaptation Evolution Strategy (CMA-ES) (Hansen and Ostermeier, 2001), improving the selection of the fittest solutions which will be perturbed to generate new elites. Mix-ME (Ingvarsson et al., 2023) extends MAP-Elites to multi-agent domains, but is limited to fully cooperative settings.
@@
Recent advancements in the field of cooperative multi-agent RL (Foerster et al., 2018; Rashid et al., 2018; de Witt et al., 2020; Mahajan et al., 2019) have shown remarkable success in tackling complex challenges in video games, such as StarCraft II (Samvelyan et al., 2019; Ellis et al., 2023). Google Research Football (GRF, Kurach et al., 2020) stands as one of the most complex multi-agent RL benchmarks, as a two-team zero-sum game with sparse reward and requiring significant amount of coordination between co-players. Most of the prior work on addresses the toy settings of the GRF only involved a few agents (e.g., 3 vs 1 scenario). Multi-Agent PPO (MAPPO, Yu et al., 2022) uses PPO (Schulman et al., 2017) with a centralised critic to play on toy settings. CDS (Li et al., 2021) analyses the importance of diversity between policies in GRF. Multi-Agent Transformer (MAT, Wen et al., 2022) models GRF as a sequence problem using the self-attention mechanism. TiKick (Huang et al., 2021) attempts to solve the full 11 vs 11 game using demonstrations from single-agent trajectories. SPC (Wang et al., 2023b) uses an adaptive curriculum on handcrafted environments for overcoming the sparse reward issue in GRF. TiZero is the first method that claims to have mastered the full 11 vs 11 game of GRF from scratch (Lin et al., 2023) following $45$ days of training with large amount of computational resources. To achieve this, TiZero uses a hand-crafted curricula over environment variations, self-play, augmented observation space, reward shaping, and action masking. Of notable importance are also the works tackling the game of football in a robotics setting (Kitano et al., 1997a, b; Riedmiller et al., 2009).
@@
Deep neural networks, such as image classifiers, are known to be sensitive to adversarial attacks (Szegedy et al., 2013; Carlini et al., 2019; Ren et al., 2020). Such susceptibility has also been demonstrated in multi-agent RL. (Wang et al., 2023a) attacks the leading Go-playing AI, KataGo (Wu, 2019), by training adversarial policies and achieving ¿97% win rate against it. Such adversarial agents are not expert Go-playing bots at all and are easily defeated by amateur human players, instead they simply trick KataGo into making serious blunders. Similarly, (Timbers et al., 2022) introduce ISMCTS-BR, a search-based deep RL algorithm that learns a best response to a given agent. Both of these solutions find exploitability using RL and expensive Monte-Carlo tree search (Coulom, 2007), whereas MADRID is a fast, gradient-free, training-free method that finds adversarial settings using QD. Furthermore, unlike the previous methods, MADRID is not restricted to any concrete agent architecture and is more general in nature. MAESTRO (Samvelyan et al., 2023) crafts adversarial curricula for training robust agents in 2-player settings by jointly sampling environment/co-player pairs, emphasizing the interplay between agents and environments.
@@ 7. Conclusion and Future Work
This paper introduced Multi-Agent Diagnostics for Robustness via Illuminated Diversity (MADRID), a novel approach aimed at systematically uncovering situations where pre-trained multi-agent RL agents display strategic errors in complex environments. MADRID leverages quality-diversity mechanisms and employs the concept of regret to identify and quantify a multitude of scenarios where agents enact suboptimal strategies, with a particular focus on the advanced TiZero agent within the Google Research Football environment. Our investigations using MADRID revealed several previously unnoticed vulnerabilities in TiZero’s strategic decision-making, such as ineffective finishing and misunderstandings of the offside rule, highlighting the hidden strategic inefficiencies and latent vulnerabilities in even the most advanced RL agents.
Looking forward, we are eager to use MADRID in broader multi-agent domains, combining it with advanced evolutionary and learning strategies to enhance its ability to pinpoint strategic inefficiencies. Investigating various adversarial situations and RL models will offer deeper understanding of strategic complexity and learning in multi-agent systems. This will aid in creating more robust RL solutions. Future research with MADRID should aim at optimising strategies to identify and correct strategic errors, furthering the advancement of robust multi-agent systems.
@@ Appendix A Adversarial Examples for Google Research Football
Below are 11 adversarial examples in TiZero we identifying using MADRID.
@@
Despite its strong performance under standard evaluations, TiZero frequently falls victim to erroneously passing the ball to players unmistakably in offside positions, as shown in Figure 12 This observations highlights TiZero’s lack of a deep understanding of the rules of the game. In contrast, the reference policies abstain from passing the ball to offside players, resulting in successful scoring outcomes. [^8] A player is offside when it is in the opponents’ half and any part of their body is closer to the opponents’ goal line than both the ball and the second-last opponent. Usually one of the two opponents is the goalkeeper. When this happens a free kick is awarded to the opponent’s team.
@@
(b) The receiving player is clearly in offside, thus a freekick is awarded to the opponents team.
(c) Reference policy does not pass to offside player and directly runs towards the goal to score.
@@
Perhaps the most glaring adversarial behaviour discovered are instances where TiZero agents inexplicably shoot towards their own goal, resulting in unforced own goals (See Figure 13). In contrast, when starting from identical in-game positions, the reference policies manage to counterattack effectively, often resulting in successful scoring endeavors.
@@
Slow running opponents
The TiZero agents always choose to sprint throughout the episode. However, this makes them weak on defense against opponents who move slower with the ball. Instead of trying to tackle and take the ball, TiZero’s main defensive strategy is to try and block opponents. Opponents can take advantage of this by using deceptive moves, especially when moving slowly, making it hard for TiZero’s defenders to stop them. This is illustrated in Figure 14.
Figure 14. Adversarial example of a slow running opponent. Three TiZero-controlled defenders are not able to stop a simple slow running opponent controlled by the reference policy, who walks past them and scores.
@@
(b) TiZero shoots from a narrow angle is blocked by the goalkeeper
@@
A notable shortcoming in TiZero’s policy, when compared to the built-in heuristic, is its reluctance to pass the ball to teammates who are in more favorable positions and have a higher likelihood of scoring, as illustrated in Figure 16. In contrast, heuristic bots—whether easy, medium, or hard—demonstrate a consistent pattern of passing to optimally positioned players, enhancing their goal-scoring opportunities. This effective passing strategy seems unfamiliar to TiZero, causing it difficulty in overcoming a successful defense.
@@
(c) Reference policy passes the ball to a better positioned player who scores.
@@
Capitalizing on another game mechanics, the reference policies exhibit stronger behaviours by halting their sprinting behaviour leading up to a shot, resulting in a notably higher success rate in goal realisation. TiZero’s agents, in contrast, consistently maintain a sprinting stance, thereby frequently missing straightforward scoring opportunities in front of the opposing goalkeepers (Figure 17).
@@
Confused Agent Behavior
@@
(c) Reference policy attacks the opponent goal, often resulting in goal scoring endeavours.
@@
TiZero shows several vulnerabilities in its defensive strategies, failing to close down on the opponent attacking trajectory and allowing them to score. In comparison, Figure 19 shows the reference policies closing down on the opponent striker and seizing the ball before they have the chance to shoot.
@@
Several adversarial examples show the entirety of TiZero’s team running in the wrong direction to defend their goal, while the ball is positioned favourably towards the opponents goal, leaving a solitary attacking player without support, who gets deceived and performs poorly. The reference policy instead doesn’t get tricked and often manages to score despite the disarray (Figure 20).
@@
TiZero often fails to acknowledge easy goal scoring opportunity, where it could get to the ball and score, but instead decides not to pursue it. Figure 22 shows how the reference policy capitalises on this kind of opportunity and scores.
@@
(b) TiZero’s attacker does not realise it can get to the ball before the goalkeeper, and runs backwards.
@@
Figure 22. Adversarial example of missing a goal scoring opportunity.

### v2

@@ 5. Results and Discussion
In our analysis of targeting TiZero on GRF, we closely examine the performance of MADRID and baselines. Figure 4(a) displays the average estimated regret values for all $160$ cells within the MAP-Elites archive across the entire collection of reference policies. Here, MADRID outperforms both baselines. The random baseline exhibits a negative value close to 00, as TiZero proves to be a stronger policy than all the reference policies on entirely random game levels. On the other hand, the targeted baseline iteratively populates the archive with high-regret levels, closely resembling MADRID’s performance at the early stages of iterations. However, as the iterations continue, it lags behind due to its failure to capitalise on previously identified high-regret levels that serve as stepping stones for the next iterations.
In Figure 4(b), we illustrate the variation in estimated final regret scores across the reference policies. Here, the regret increases as we move to higher-ranked agents. The heuristic bots display regret levels that are on par with the intermediate checkpoints of TiZero.
As we approximate the regret using the difference between cross-play (XP) and self-play (SP) between reference and TiZero policies (see Equations 1 and 2), a regret estimate of $1$ for an adversarial level $\theta$ can be achieved in two situations. First, the reference policy scores against TiZero in XP, while TiZero cannot score in SP. Second, TiZero concedes a goal in SP in $\theta$. Intriguingly, our findings reveal that for around 90% of the adversarial levels generated by MADRID, a nominally weaker reference policy outperforms TiZero. This emphasises MADRID’s capability in exposing adversarial levels where even state-of-the-art policies are prone to missteps.
@@
Figure 4(c) and Figure 4(d) illustrate the estimated rate of goals scored against TiZero by the reference policies on adversarial levels produced by MADRID and baselines. We can see that approximately 70% of the time across all reference policies, the reference policy scored a goal against TiZero in a short period of time. [^6] The levels last only $128$ environment steps, which is a short episode compared to the $3000$ steps for the full game. It should be noted that within the remaining 30%, the majority of instances result in no scored goals.
Figure 5 highlights the difference in performance for selected reference policies. Notably, the higher-ranked checkpoints of TiZero, saved at the later stages of its training, can be used to identify more severe vulnerabilities, as measured using the regret estimate.
Figure 6 shows the evolution of MADRID’s archive for a specific reference policy, illustrating its search process over time. Initially, the grid is sparsely filled with low-regret levels. However, as iterations progress, MADRID generates high-regret levels that progressively populate the entire grid. This shows that MADRID can discover high-regret levels anywhere on the football field. On average, we notice that higher-level scenarios tend to be located towards the positive $x$ coordinates. These correspond to situations where the ball is close to the opponent’s goal from the perspective of the reference policy. While most regret scores tend to have uniform values around similar positions on the field, in Figure 6(d) the grid also includes an adversarial level with an estimated regret of $1.75$. This indicates that MADRID found a level where the reference policy scores against TiZero in XP, while TiZero concedes a goal in SP.
@@ 5.1. Qualitative Analysis
We conduct a qualitative analysis of the adversarial levels identified by MADRID on GRF by visualising the highest ranking levels in the archive across all reference policies. We provide a selection of these examples below, with a comprehensive list available in Appendix A. Full videos of all identified vulnerabilities can be found at https://sites.google.com/view/madrid-marl.
@@
(b) The receiving player is clearly offside, thus a freekick is awarded to the opponent team.
(c) Reference policy does not pass to the offside player and directly runs towards the goal to score.
@@
Despite its strong performance under standard evaluations, TiZero frequently falls victim to erroneously passing the ball to players unmistakably in offside positions, as shown in Figure 7 This observation highlights TiZero’s lack of a deep understanding of the rules of the game. In contrast, the reference policies abstain from passing the ball to offside players, resulting in successful scoring outcomes. [^7] Player are offside when they are in the opponent’s half and any part of their body is closer to the opponent’s goal line than both the ball and the second-last opponent. Usually one of the two opponents is the goalkeeper. When the ball is passed to a player who is offside, a free kick is awarded to the opponent’s team.
@@
Figure 9. Adversarial example of a slow-running opponent. Three TiZero-controlled defenders are not able to stop a simple slow-running opponent, who walks past them and scores.
@@
(c) Reference policy passes the ball to a better-positioned player who scores.
@@
Perhaps the most glaring adversarial behaviour discovered are instances where TiZero agents inexplicably shoot towards their own goal, resulting in unforced own goals (See Figure 8). In contrast, when starting from identical in-game positions, the reference policies manage to counterattack effectively, often resulting in successful scoring endeavours.
Slow-running opponents
The TiZero agents always choose to sprint throughout the episode. However, this makes them weak on defence against opponents who move slower with the ball. Instead of trying to tackle and take the ball, TiZero’s main defensive strategy is to try and block opponents. Opponents can take advantage of this by using deceptive moves, especially when moving slowly, making it hard for TiZero’s defenders to stop them. This is illustrated in Figure 9.
@@
A notable shortcoming in TiZero’s policy, when compared to the built-in heuristic, is its reluctance to pass the ball to teammates who are in more favourable positions and have a higher likelihood of scoring, as illustrated in Figure 11. In contrast, heuristic bots—whether easy, medium, or hard—demonstrate a consistent pattern of passing to optimally positioned players, enhancing their goal-scoring opportunities. This effective passing strategy seems unfamiliar to TiZero, causing it difficulty in overcoming a successful defence.
@@ 6. Related Work
Quality Diversity (QD) is a category of open-ended learning methods aimed at discovering a collection of solutions that are both highly diverse and performant (Lehman and Stanley, 2011; Cully and Demiris, 2018). Two commonly used QD algorithms are Novelty Search with Local Competition (NSLC, Lehman and Stanley, 2011) and MAP-Elites (Mouret and Clune, 2015; Cully et al., 2015). These two approaches differ in the way they structure the archive; novelty search completely forgoes a grid and opts instead for growing an unstructured archive that dynamically expands, while MAP-Elites adopts a static mapping approach. Although MADRID leverages MAP-Elites as its diversity mechanism, it can be adapted to use NSLC. One of the most effective versions of MAP-Elites is CMA-ME (Fontaine et al., 2020). CMA-ME combines MAP-Elites with the evolutionary optimisation algorithm Covariance Matrix Adaptation Evolution Strategy (CMA-ES) (Hansen and Ostermeier, 2001), improving the selection of the fittest solutions which will be perturbed to generate new elites. Mix-ME (Ingvarsson et al., 2023) extends MAP-Elites to multi-agent domains but is limited to fully cooperative settings.
@@
Recent advancements in the field of cooperative multi-agent RL (Foerster et al., 2018; Rashid et al., 2018; de Witt et al., 2020; Mahajan et al., 2019) have shown remarkable success in tackling complex challenges in video games, such as StarCraft II (Samvelyan et al., 2019; Ellis et al., 2023). Google Research Football (GRF, Kurach et al., 2020) stands as one of the most complex multi-agent RL benchmarks, as a two-team zero-sum game with sparse reward and requiring a significant amount of coordination between co-players. Most of the prior work on addressing the toy settings of the GRF only involved a few agents (e.g., 3 vs 1 scenario). Multi-Agent PPO (MAPPO, Yu et al., 2022) uses PPO (Schulman et al., 2017) with a centralised critic to play on toy settings. CDS (Li et al., 2021) analyses the importance of diversity between policies in GRF. Multi-Agent Transformer (MAT, Wen et al., 2022) models GRF as a sequence problem using the self-attention mechanism. TiKick (Huang et al., 2021) attempts to solve the full 11 vs 11 game using demonstrations from single-agent trajectories. SPC (Wang et al., 2023b) uses an adaptive curriculum on handcrafted environments for overcoming the sparse reward issue in GRF. TiZero is the first method that claims to have mastered the full 11 vs 11 game of GRF from scratch (Lin et al., 2023) following $45$ days of training with a large amount of computational resources. To achieve this, TiZero uses a hand-crafted curriculum over environment variations, self-play, augmented observation space, reward shaping, and action masking. Of notable importance are also the works tackling the game of football in a robotics setting (Kitano et al., 1997a, b; Riedmiller et al., 2009).
@@
Deep neural networks, such as image classifiers, are known to be sensitive to adversarial attacks (Szegedy et al., 2013; Carlini et al., 2019; Ren et al., 2020). Such susceptibility has also been demonstrated in multi-agent RL. Wang et al. (2023a) attacks the leading Go-playing AI, KataGo (Wu, 2019), by training adversarial policies and achieving ¿97% win rate against it. Such adversarial agents are not expert Go-playing bots and are easily defeated by amateur human players. Instead, they simply trick KataGo into making serious blunders. Timbers et al. (2022) introduce ISMCTS-BR, a search-based deep RL algorithm that learns a best response to a given agent. Both of these solutions find exploitability using RL and expensive Monte-Carlo tree search (Coulom, 2007), whereas MADRID is a fast, gradient-free, black-box method that finds adversarial settings using QD. Unlike the previous methods, MADRID is not restricted to any concrete agent architecture and is more general in nature. MAESTRO (Samvelyan et al., 2023) crafts adversarial curricula for training robust agents in 2-player settings by jointly sampling environment/co-player pairs, emphasizing the interplay between agents and environments.
@@ 7. Conclusion and Future Work
This paper introduced Multi-Agent Diagnostics for Robustness via Illuminated Diversity (MADRID), a novel approach aimed at systematically uncovering situations where pre-trained multi-agent RL agents display strategic errors. MADRID leverages quality-diversity mechanisms and employs regret to identify and quantify a multitude of scenarios where agents enact suboptimal strategies, with a particular focus on the advanced TiZero agent within the Google Research Football environment. Our investigations using MADRID revealed several previously unnoticed vulnerabilities in TiZero’s strategic decision-making, such as ineffective finishing and misunderstandings of the offside rule, highlighting the hidden strategic inefficiencies and latent vulnerabilities in even the most advanced RL agents.
Looking forward, we are eager to use MADRID in broader multi-agent domains, combining it with advanced evolutionary and learning strategies to enhance its ability to pinpoint limitations of multi-agent policies. Investigating various adversarial situations in pre-trained models will offer a deeper understanding of strategic complexities within multi-agent systems. Future research will also aim at using the discovered vulnerabilities in pre-trained models for enhancing the target model further through additional fine-tuning, thereby improving its robustness to unseen or adversarial settings.
@@ Appendix A Adversarial Examples for Google Research Football
Below are 11 adversarial examples in TiZero we identify using MADRID.
@@
Despite its strong performance under standard evaluations, TiZero frequently falls victim to erroneously passing the ball to players unmistakably in offside positions, as shown in Figure 12 This observation highlights TiZero’s lack of a deep understanding of the rules of the game. In contrast, the reference policies abstain from passing the ball to offside players, resulting in successful scoring outcomes. [^8] A player is offside when it is in the opponent’s half and any part of their body is closer to the opponent’s goal line than both the ball and the second-last opponent. Usually one of the two opponents is the goalkeeper. When this happens a free kick is awarded to the opponent’s team.
@@
(b) The receiving player is clearly offside, thus a freekick is awarded to the opponent’s team.
(c) Reference policy does not pass to the offside player and directly runs towards the goal to score.
@@
Perhaps the most glaring adversarial behaviour discovered are instances where TiZero agents inexplicably shoot towards their own goal, resulting in unforced own goals (See Figure 13). In contrast, when starting from identical in-game positions, the reference policies manage to counterattack effectively, often resulting in successful scoring endeavours.
@@
Slow-running opponents
The TiZero agents always choose to sprint throughout the episode. However, this makes them weak in defence against opponents who move slower with the ball. Instead of trying to tackle and take the ball, TiZero’s main defensive strategy is to try and block opponents. Opponents can take advantage of this by using deceptive moves, especially when moving slowly, making it hard for TiZero’s defenders to stop them. This is illustrated in Figure 14.
Figure 14. Adversarial example of a slow-running opponent. Three TiZero-controlled defenders are not able to stop a simple slow-running opponent controlled by the reference policy, who walks past them and scores.
@@
(b) TiZero shoots from a narrow angle and is blocked by the goalkeeper
@@
A notable shortcoming in TiZero’s policy, when compared to the built-in heuristic, is its reluctance to pass the ball to teammates who are in more favourable positions and have a higher likelihood of scoring, as illustrated in Figure 16. In contrast, heuristic bots—whether easy, medium, or hard—demonstrate a consistent pattern of passing to optimally positioned players, enhancing their goal-scoring opportunities. This effective passing strategy seems unfamiliar to TiZero, causing it difficulty in overcoming a successful defence.
@@
(c) Reference policy passes the ball to a better-positioned player who scores.
@@
Capitalizing on other game mechanics, the reference policies exhibit stronger behaviours by halting their sprinting behaviour leading up to a shot, resulting in a notably higher success rate in goal realisation. TiZero’s agents, in contrast, consistently maintain a sprinting stance, thereby frequently missing straightforward scoring opportunities in front of the opposing goalkeepers (Figure 17).
@@
Confused Agent Behaviour
@@
(c) Reference policy attacks the opponent’s goal, often resulting in goal-scoring endeavours.
@@
TiZero shows several vulnerabilities in its defensive strategies, failing to close down on the opponent’s attacking trajectory and allowing them to score. In comparison, Figure 19 shows the reference policies closing down on the opponent striker and seizing the ball before they have the chance to shoot.
@@
Several adversarial examples show the entirety of TiZero’s team running in the wrong direction to defend their goal, while the ball is positioned favourably towards the opponent’s goal, leaving a solitary attacking player without support, who gets deceived and performs poorly. The reference policy instead doesn’t get tricked and often manages to score despite the disarray (Figure 20).
@@
TiZero often fails to acknowledge easy goal-scoring opportunities, where it could get to the ball and score, but instead decides not to pursue it. Figure 22 shows how the reference policy capitalises on this kind of opportunity and scores.
@@
(b) TiZero’s attacker does not realise it can get to the ball before the goalkeeper and runs backwards.
@@
Figure 22. Adversarial example of missing a goal-scoring opportunity.